OpenStudy (anonymous):

Solve 4r2 – 9 = 91. A. {± 3} B. {± 4} C. {± 5} D. {±6}

OpenStudy (anonymous):

4r^2-9=91?

OpenStudy (anonymous):

add +9 for both sides 4r^2-9 +9=91+9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4r^2=100 r^2=25 r=+-5

OpenStudy (kendricklamar2014):

1. A 9 to both sides:\[4r^2−9+9=91+9 = 4r^2=100\]

OpenStudy (anonymous):

so c? @ssssss

OpenStudy (kendricklamar2014):

2. Divide both sides by 4:\[\frac{ 4r^2 }{ 4 } = \frac{ 100 }{ 4 }\]

OpenStudy (anonymous):

Yes))

OpenStudy (kendricklamar2014):

Yes, C is Correct :)

OpenStudy (anonymous):

you r left with r2 and 25 @KendrickLamar2014

OpenStudy (kendricklamar2014):

Yes, then u take the square root of 25

OpenStudy (anonymous):

r^2=100/4=25 r=+-5

OpenStudy (anonymous):

is 5 so c?

OpenStudy (kendricklamar2014):

Which equals: r = 5 or r = -5

OpenStudy (kendricklamar2014):

C is o r r e c t

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Solve (z + 6)2 = 5. A. B. C. D.

OpenStudy (kendricklamar2014):

One question per post